OpenStudy (anonymous):

Find the Taylor series of fx = cosx around a= pi. Find it's radius of convergence

OpenStudy (anonymous):

I took this today lol >_< let me see what I can do:\[\cos x = \sum_{k=0}^{\infty} \frac{(-1)^k}{(2k)!} (x- \pi)^{2k}\] hmm, I didn't study it yet, but I have tried lol, I'm sorry ^_^"

OpenStudy (anonymous):

when c = pi

OpenStudy (anonymous):

but if c = 0, then (x)^2k